23. Flies with red eyes are homozygous (RR) for the reddish eye pigment.

Flies with orange eyes are heterozygous (Rr), and white-eyed flies are homozygous recessive (rr).

You have a population with the following frequencies

36 red eyes + 48 orange + 16 white eyes.

i. Determine the allele frequencies for R and r.

R =

r =

ii. What are the phenotypic frequencies predicted by Hardy-Weinberg?


iii. Is this population at Hardy-Weinberg equilibrium?
Answer #1

i)

p = 2NRR + NRr   / 2N

= 72 + 48 / 200 = 0.6

q = 2Naa + NAa  / 2N

= 32 + 48 / 200 = 0.4

Hence, p = 0.6 , and q = 0.4

R = 0.6 , r = 0.4

This is in hardy weinberg equilibrium, p + q = 0.6 + 0.4 = 1

ii) Phenotypic frequencies

RR = 36/ 100 = 0.36

Rr = 48 /100 = 0.48

rr = 16/100 = 0.16

iii) Yes the population is in hardy weinberg equilibrium , p + q= 1